OpenStudy (anonymous):

9-x=16 wouldn't you subtract 9 from both sides?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

9-x = 16 -x = 7 x = -7

OpenStudy (lalaly):

yes -x=7 x=-7

OpenStudy (anonymous):

x =-7

OpenStudy (anonymous):

that's what thought ;)